OpenStudy (anonymous):

The difference of the areas of two circles is 21π. If their radii are r+3 and r, find the radius of the larger circle.

OpenStudy (anonymous):

5

OpenStudy (anonymous):

how u get?

OpenStudy (anonymous):

pi(r+3)^2 - pi(r)^2 = 21 pi

myininaya (myininaya):

r^2+6r+9-r^2=21 6r=21-9 6r=12 r=2 r+3=2+3=5 gj fax!